Can you pleass show the work and explain?

(1) How many grams of potassium chloride must be dissolved in 2.50 x 10^(2) mL of water to make a 0.400 m solution?

(2) How many milliliters of commercial hydrochloric acid, which is 11.6 molar, should you use to prepare 5.50 L of 0.500 molar hydrochloric acid?

(3) Calculate the molarity of a NaOH solution prepared by mixing 100. mL of 0.20 M NaOH with 150 mL of H2O.


(4) Calculate the volume in milliliters of 0.842 M sodium hydroxide required to precipitate as copper(II) hydroxide all of the copper ions in 30.0 mL of 0.635 M copper(II) sulfate
Answer #1

1) For water density = 1g/ml

As density = mass/volume

So, 1 = mass/volume

Hence mass = volume = 2.50×102 g

Mass of KCl = Molality×molar mass×mass of water/1000

= 0.400×74.55×2.50×102/1000

= 7.455 g

2) V1 = M2V2/M1 = 0.500×5500/11.6 = 237.06 ml

3) M2 = M1V1/V2 = 0.2×100/150 = 0.13 M

4) Moles of copper = molarity×volume/1000

= 0.635×30/1000 = 0.01905 moles

Moles of NaOH = 2×0.01905 = 0.0381 moles

Volume of NaOH = moles×1000/molarity

= 0.0381×1000/0.842 = 45.25 ml